# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 2.1 Problem Statement
A simple situation where Assumption 1 holds when $X^{\mathrm{in}}$ is a bounded set, $T$ maps from bounded sets to bounded sets and $\varphi$ is upper semicontinuous.
2.2 An unstable linear system and an objective function with finite optimal value
Throughout the paper, we illustrate our method and the computations on the example presented in this subsection. The peaks computation problem can be solved more simply and we compare the results obtained with our methodology with the one obtained analytically.
Let us consider the following dynamics on $\mathbb{R}^{2}$:
$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}.$
We parameterize the initial conditions set by a scalar $\mu\in(0,1/3]$ and we define:
$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}.$
Finally, the objective function is parameterized by a positive scalar $p$:
$\mathbb{R}^{2}\ni(x,y):\mapsto\varphi(x,y):=y^{2}-x^{2}+px$
The associated peaks computation problem is the following:
$\begin{array}[]{rc}\displaystyle{\operatorname*{Max}_{(x,y),n}}&\varphi(A^{n}(x,y)^{\intercal})\\ s.t.&\left\{\begin{array}[]{l}(x,y)\in X^{\mathrm{in}}\\ n\in\mathbb{N}\end{array}\right.\end{array}$
The system is linear and the objective function is quadratic. However, we remark that $A$ admits $1/2$ and $3/2$ as eigenvalues. Therefore, $A$ is unstable and for some $x$, $\|A^{n}x\|$ goes to +$\infty$+ ∞ as $n$ goes to +$\infty$+ ∞. Therefore, we cannot use the method developed in [Adj23]. Indeed, the system does not admit a quadratic Lyapunov function. We also note that $\varphi$ is not concave nor convex.
As the matrix $A$ is diagonalizable, the powers of it are easy to compute and for all $n\in\mathbb{N}$ we have :
$A^{n}=\dfrac{1}{2^{n+1}}\begin{pmatrix}3^{n}+1&2(3^{n}-1)\\ \frac{3^{n}-1}{2}&3^{n}+1\end{pmatrix}\text{ and }\forall\,y\in\mathbb{R},\ A^{n}\begin{pmatrix}2y\\ y\end{pmatrix}=\left(\frac{3}{2}\right)^{n}\begin{pmatrix}2y\\ y\end{pmatrix}$
Moreover, the initial conditions set is actually a segment each element of it can be written as $(2y,y)^{\intercal}$ with $y\in[\mu,1/2]$. Using this, we have for all $y\in[\mu,1/2]$, for all $n\in\mathbb{N}$:
$\varphi_{n}(y):=\varphi(A^{n}(2y,y)^{\intercal})=-3\left(\dfrac{3}{2}\right)^{2n}y^{2}+2p\left(\dfrac{3}{2}\right)^{n}y$
Let $n\in\mathbb{N}$. The function $y\mapsto\varphi_{n}(y)$ is concave and infinitely derivable. It admits an unique maximizer on $\mathbb{R}$ which is $y_{n}^{*}:=\left(\frac{2}{3}\right)^{n}\frac{p}{3}$. The optimal value of $\varphi_{n}$ on $\mathbb{R}$ is $p^{2}/3$. The function $\varphi_{n}$ is strictly increasing before $y_{n}^{*}$ and strictly decreasing after. As we maximize on $[\mu,1/2]$, we have to check whether $y_{n}^{*}$ belongs to this interval. To avoid the case where $\frac{1}{2}$ is not used as a maximizer we assume that $p>3/2$. In this case, $y_{0}^{*}>1/2$.
As the sequence $(y_{n}^{*})_{n\in\mathbb{N}}$ strictly decreases to zero, we can define two integers:
$\underline{n}:=\lfloor\ln(3/2p)/\ln(2/3)\rfloor+1\text{ and }\overline{n}:=\lfloor\ln(3\mu/p)/\ln(2/3)\rfloor+1$
The integer $\underline{n}$ corresponds to the smallest integer such that $y_{n}^{*}<1/2$ whereas $\overline{n}$ is the smallest integer such that $y_{n}^{*}<\mu$. Moreover, as we assume that $\mu\in(0,1/3]$, we have $(\ln(3\mu/p)-\ln(3/2p))/\ln(2/3)=-(\ln(\mu)+\ln(2))\ln(3/2)\geq-(\ln(1/3)+\ln(2))/\ln(3/2)=1$. We conclude that $\overline{n}>\underline{n}$, as for all $x,y\in\mathbb{R}_{+}$ such that $x-y\geq 1$, we have $\lfloor x\rfloor\geq\lfloor y\rfloor+1$.
Finally using $\underline{n}$ and $\overline{n}$ , we can define the sequence of interest $(\nu_{n})_{n\in\mathbb{N}}$ and for all $n\in\mathbb{N}$:
$\nu_{n}=\left\{\begin{array}[]{lr}\varphi_{n}(1/2)&\text{ if }n\in[0,\underline{n}-1]\cap\mathbb{N}\\ \dfrac{p^{2}}{3}&\text{ if }n\in[\underline{n},\overline{n}-1]\cap\mathbb{N}\\ \varphi_{n}(\mu)&\text{ if }n\geq\overline{n}\end{array}\right.$ (2)
From now on, when we refer to this example, we will assume that $\mu\in(0,1/3]$ and $p\geq 3/2$.
3 Suprema and Rank Maximizer of General Sequences
3.1 Notations and Basic Results
We introduce the set of real sequences which are bounded from above:
$\Lambda:=\{u=(u_{0},u_{1},\ldots)\in\mathbb{R}^{\mathbb{N}}:\sup_{n\in\mathbb{N}}u_{n}<+\infty\}.$
For an element of $\Lambda$, we are interested in computing the supremum of its terms.
@@ 3.1 Notations and Basic Results
$u\in\Lambda\iff\displaystyle{\limsup_{n\to+\infty}u_{n}\in\mathbb{R}\cup\{-\infty\}}$.
For $u\in\Lambda$, we are also looking for the set of maximizers i.e. $\operatorname{Argmax}(u):=\{k\in\mathbb{N}:u_{k}=\displaystyle{\sup_{n\in\mathbb{N}}u_{n}}\}$. For computations purpose, we need to characterize $\operatorname{Argmax}(u)$. Consequently, we introduce, for $u\in\Lambda$, the two sets of ranks:
$\Delta_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}\right\}\ \text{ and }\ \Delta^{s}_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}>\sup_{j>k}u_{j}\right\}$
It is easy to see that $\Delta_{u}\subseteq\Delta^{s}_{u}$. Moreover, if $k$ belongs to $\Delta_{u}$ (resp. $\Delta^{s}_{u}$), then any integer greater than $k$ belongs to $\Delta_{u}$ (resp. $\Delta^{s}_{u}$). Besides, we will need the set of ranks for which the associated term is greater than the limsup (resp. strictly greater):
${\mathrm{R}}_{u}:=\left\{k\in\mathbb{N}:u_{k}\geq\limsup_{n\to+\infty}u_{n}\right\}\ \text{ and }\ {\mathrm{R}}_{u}^{s}:=\left\{k\in\mathbb{N}:u_{k}>\limsup_{n\to+\infty}u_{n}\right\}$
It is obvious that ${\mathrm{R}}_{u}^{s}\subseteq{\mathrm{R}}_{u}$. Note that even if $u\in\Lambda$, ${\mathrm{R}}_{u}$ and hence ${\mathrm{R}}_{u}^{s}$ can be empty. To construct the characterizations, we need to consider the smallest elements of $\Delta_{u}$, $\Delta^{s}_{u}$, ${\mathrm{R}}_{u}$ and ${\mathrm{R}}_{u}^{s}$.
$\mathbf{K}_{u}=\inf\Delta_{u};\ \qquad\mathbf{K}^{s}_{u}=\inf\Delta^{s}_{u};\ \qquad k_{u}=\inf{\mathrm{R}}_{u}\quad\text{ and }\ k^{s}_{u}=\inf{\mathrm{R}}_{u}^{s}$
We recall that the infimum of the empty set is equal to +$\infty$+ ∞. Hence, $\mathbf{K}_{u}<+\infty$ (resp. $\mathbf{K}^{s}_{u}<+\infty$) if and only if $\Delta_{u}\neq\emptyset$ (resp. $\Delta^{s}_{u}\neq\emptyset$); and $k_{u}<+\infty$ (resp. $k^{s}_{u}<+\infty$) if and only if ${\mathrm{R}}_{u}\neq\emptyset$ (resp. ${\mathrm{R}}_{u}^{s}\neq\emptyset$). Furthermore, if the sets of integers defined above are nonempty, they contain their infimum i.e. $\mathbf{K}_{u}\in\Delta_{u}$, $\mathbf{K}^{s}_{u}\in\Delta^{s}_{u}$, $k_{u}\in{\mathrm{R}}_{u}$ and $k^{s}_{u}\in{\mathrm{R}}_{u}^{s}$ when those integers are finite.
Proposition 2.
If $u\notin\Lambda$ then $\Delta_{u}=\emptyset$ and ${\mathrm{R}}_{u}=\emptyset$. Obviously, it also implies that $\Delta^{s}_{u}=\emptyset$ and ${\mathrm{R}}_{u}^{s}=\emptyset$.
Proposition 3.
Let $u\in\Lambda$. The following assertions hold:
@@
    $\Delta^{s}_{u}\neq\emptyset\iff{\mathrm{R}}_{u}^{s}\neq\emptyset$;
@@
    $\Delta_{u}\neq\emptyset\iff{\mathrm{R}}_{u}\neq\emptyset$;
@@
    $\Delta^{s}_{u}=\emptyset\iff\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$.
Proposition 4.
The following inequalities hold:
$k_{u}\leq k^{s}_{u};\qquad k_{u}\leq\mathbf{K}_{u};\qquad k^{s}_{u}\leq\mathbf{K}^{s}_{u}\quad\text{ and }\mathbf{K}_{u}\leq\mathbf{K}^{s}_{u}\enspace.$
Moreover, if $k^{s}_{u}<+\infty$, then $k^{s}_{u}\leq\mathbf{K}_{u}$.
Proposition 5.
Let $u\in\Lambda$. The following assertions hold:
@@
    $\operatorname{Argmax}(u)\subseteq\Delta_{u}$;
@@
    $\operatorname{Argmax}(u)\neq\emptyset\iff\Delta_{u}\neq\emptyset$;
  3. 3.
    $\mathbf{K}_{u}=\inf\operatorname{Argmax}(u)$;
  4. 4.
    $\operatorname{Argmax}(u)=\emptyset$ or $\operatorname{Argmax}(u)$ nonempty and unbounded $\implies\sup_{n\in\mathbb{N}}u_{n}=\limsup_{k\to+\infty}u_{k}$;
  5. 5.
    If $\Delta^{s}_{u}\neq\emptyset$ then $\mathbf{K}^{s}_{u}=\max\operatorname{Argmax}(u)$.
Note that $\operatorname{Argmax}(u)$ can be nonempty and bounded when $\Delta^{s}_{u}$ is empty. In this case, by the third statement of Prop 3, we have $\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$. However, the sequence finitely touches $\limsup_{n\to+\infty}u_{n}$ before approaching, up to a subsequence, $\limsup_{n\to+\infty}u_{n}$ from below (strictly).
Example 1.
Let us consider the peaks computation problem depicted at Subsection 2.2. The sequence $(\nu_{k})_{k\in\mathbb{N}}$ presented at Eq. (2) has a greatest maximizer as $\limsup_{k\to+\infty}\nu_{k}=\lim_{k\to+\infty}\varphi_{k}(\mu)=-\infty$. The set $\operatorname{Argmax}(u)$ would not be bounded from above if $\mu=0$ and thus $\overline{n}=+\infty$. Indeed, in this case, we would get $\limsup_{k\to+\infty}\nu_{k}=p^{2}/3=\sup_{k\in\mathbb{N}}\nu_{k}$. ∎
Proposition 6.
Let $u\in\Lambda$. The following assertions hold:
@@
    Assume that ${\mathrm{R}}_{u}\neq\emptyset$. Let $k\geq k_{u}$ such that $\sup_{k_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
@@
    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $k\geq k^{s}_{u}$ such that $\sup_{k^{s}_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
  3. 3.
    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. For all $\mathbf{K}_{u}\leq k$,
    $\sup_{0\leq j\leq k}u_{j}=\sup_{k^{s}_{u}\leq j\leq k}u_{j}=\sup_{k_{u}\leq j\leq k}u_{j}=\sup_{j\in\mathbb{N}}u_{j}\enspace.$
In summary, for $u\in\Lambda$, to compute $\sup_{k\in\mathbb{N}}u_{k}$, we need to study first the emptiness of ${\mathrm{R}}_{u}^{s}$. Indeed, if ${\mathrm{R}}_{u}^{s}$ is empty, we know that (the third statement of Prop. 3) $\sup_{n\in\mathbb{N}}u_{n}$ is equal to $\limsup_{n}u_{n}$. If ${\mathrm{R}}_{u}^{s}$ is not empty, then we have to compute $\mathbf{K}_{u}$ to find $\sup_{k\in\mathbb{N}}u_{k}$ and a maximizer (the second statement of Prop. 5). However, to identify $\mathbf{K}_{u}$, we need knowledge on the past and the future of the sequence. The good point is that any over-approximation $k$ of $\mathbf{K}_{u}$ permits to know $\sup_{k\in\mathbb{N}}u_{k}$ by computing $\sup_{0\leq j\leq k}u_{j}$ (the third statement of Prop 6).
3.2 A Formula Based on Pairs of Strictly Increasing Continuous Functions and Convergent Geometric Sequences
Let us write $\mathbb{I}\left(\mathbb{R}\right)$ the set of all intervals (convex sets) on $\mathbb{R}$ (open, closed, or semi-open). For a function $f$ from $\mathbb{R}$ to $\mathbb{R}$, we define the set of intervals:
$\mathrm{SIC}(f)=\{I\in\mathbb{I}\left(\mathbb{R}\right):f:I\mapsto f(I)\text{ is strictly increasing and continuous }\}.$
Recall that a strictly increasing function $f$ is continuous on an interval $[a,b]$ if and only if $f$ is surjective from $[a;b]$ to $[f(a);f(b)]$. Hence, if $I$ belongs to $\mathrm{SIC}(f)$, then $f$ is bijective from $I$ to $f(I)$ and we can define its inverse on $I$ i.e. a function $g:f(I)\mapsto I$ such that $f\circ g=g\circ f=\mathrm{Id}$ where $\mathrm{Id}:x\mapsto x$ (viewed as a function from $I$ to itself). For $I\in\mathrm{SIC}(f)$, we denote by ${{f}_{\left(I\right)}^{-1}}$ the inverse of $f$ on $I$. Note that ${{f}_{\left(I\right)}^{-1}}$ is also strictly increasing and continuous (surjective) from $f(I)$ to $I$. Finally, we pay a particular attention to functions strictly increasing and continuous on $[0,1]$ and we write :
The functions of $\Omega([0,1])$ are part of the upper bound on the sequence of interest $u$. To complete the upper bound, we need a convergent geometric sequence. We introduce, for $u\in\mathbb{R}^{\mathbb{N}}$ the following sets:
$\mathbb{\Gamma}(u):=\{(h,\beta)\in\Omega([0,1])\times(0,1):u_{k}\leq h(\beta^{k}),\ \forall\,k\in\mathbb{N}\}$

### v4

@@ 2.1.2 A problem on integers
A simple situation where Assumption 2 holds is when $X^{\mathrm{in}}$ is a bounded set, $T$ maps bounded sets to bounded sets, and $\varphi$ is upper semicontinuous. The existence of an optimal solution for Problem ($P_{k}$) is actually not mandatory to ensure Assumption 2.
From these notations, $\nu_{\rm opt}$ is equal to the optimal value of Problem (DPCP), whereas $\nu_{k}$ is the optimal value of static optimization Problem  ($P_{k}$). Solving Problem (DPCP) involves computing $\nu_{\rm opt}$, an integer $\overline{k}$ such that $\nu_{\overline{k}}=\nu_{\rm opt}$ and an optimal solution $x\in X^{\mathrm{in}}$ of Problem ($P_{k}$). In this paper, we focus on the characterization and computation of the integer $\overline{k}$. Static problems of the form ($P_{k}$) are standard problems in optimization theory. The computation of an optimal solution for these problems relies on an optimization solver when a suitable solver is available. In theory, some special instances of Problem ($P_{k}$) can be solved numerically, more or less efficiently. For example, when $\varphi\circ T^{k}$ is concave and $X^{\mathrm{in}}$ is convex, numerous methods exist (see e.g.,  [BTN01, Ber15]), or when $\varphi\circ T^{k}$ is polynomial and $X^{\mathrm{in}}$ is a basic semialgebraic set (see e.g.,  [AL11, Las15]). In practice, the difficulty of resolving Problem ($P_{k}$) precedes the optimization procedure and remains in the writing of the optimization problem itself. Indeed, obtaining a closed form for $T^{k}$ or at least for $T^{k}(X^{\mathrm{in}})$ can be an ambitious task. For example, for logistic sequences, closed forms are only known for particular cases (see, e.g.,[GRR10]). Some specific difference equations or systems require intensive efforts to obtain closed forms for $T^{k}$ (see, e.g., [Els12, HR18, Ste14] and references therein). Similarly, obtaining an exact description of $T^{k}(X^{\mathrm{in}})$ when $X^{\mathrm{in}}$ is an infinite set incurs a high computational cost. For example, in [MHL15], the authors approximate the polynomial image of a basic semialgebraic set obtained from the $L^{1}$ limit of a sequence of semidefinite programs. In some specific cases, $T^{k}$ can be described explicitly. When $T$ is linear or conjugate to a linear map (i.e., $T=B^{-1}\circ A\circ B$ for a bijection $B$ and a linear map $A$), we can compute $T^{k}$ using an eigendecomposition if $T$ is diagonalizable (or in Jordan form). When $T$ is linear, $T^{k}(X^{\mathrm{in}})$ is always a polytope if $X^{\mathrm{in}}$ is so. Hence, when $T$ is linear and $X^{\mathrm{in}}$ is a polytope, Problem ($P_{k}$) becomes a linear program. While the theory developed in this paper is independent of the practical constructions of $T^{k}$ or $T^{k}(X^{\mathrm{in}})$, the computational illustration proposed in this paper is limited to an unstable linear system combined with a nonlinear objective function $\varphi$. A more general practical approach will be considered in future works.
The approach developed in [Adj25b] consists in computing an upper bound of the greatest integer maximizer of Problem (DPCP). In [Adj25a], we employed the notion stopping integer, which is an integer $K$ such that
$\nu_{\rm opt}=\max\{\nu_{k}:k=0,\ldots,K\}\text{ and }\sup_{k>K}\nu_{k}<\nu_{\rm opt}.$ (3)
We define the set of stopping integers and the smallest stopping integer as follows:
$\Delta^{s}_{\nu}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}\nu_{j}>\sup_{j>k}\nu_{j}\right\}\text{ and }\mathbf{K}^{s}_{u}:=\inf\Delta^{s}_{\nu}.$
The set of stopping integers can be empty, and we can characterize its nonemptiness from the analysis of the limit superior of $\nu$ and thus introduce the set:
${\mathrm{R}}_{\nu}^{s}:=\left\{k\in\mathbb{N}:\nu_{k}>\limsup_{n\to+\infty}\nu_{n}\right\}.$
Moreover, we define the set of integer maximizers of Problem (DPCP) as
$\operatorname{Argmax}(\nu):=\{k\in\mathbb{N}:\nu_{k}=\nu_{\rm opt}\}.$
Recall that $\nu_{\rm opt}$ is finite if and only if the limit superior of $\nu$ is finite or equal to -$\infty$. This basic analysis can be completed using the sets $\Delta^{s}_{\nu}$ and ${\mathrm{R}}_{\nu}^{s}$.
@@
If $\nu_{\rm opt}<+\infty$. The following assertions are made:
@@
    $\Delta^{s}_{\nu}\neq\emptyset\iff{\mathrm{R}}_{\nu}^{s}\neq\emptyset$;
@@
    $\Delta^{s}_{\nu}=\emptyset\iff\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$;
@@
    If $\Delta^{s}_{\nu}\neq\emptyset$ then $\mathbf{K}^{s}_{\nu}=\max\operatorname{Argmax}(\nu)$.
One way to compute an element of ${\mathrm{R}}_{\nu}^{s}$ is, given a term $\nu_{k}$, to look for an integer $K\in\mathbb{N}$ such that
$\nu_{j}<\nu_{k},\text{ for all }j\geq K.$
Such integers $K$ cannot be smaller than $\mathbf{K}^{s}_{\nu}$ and are thus stopping integers. The problem is to convert the term $\nu_{k}$ into a stopping integer. The most natural sequences allowing this conversion (which is a bijection) between terms and indices are the geometric sequences. However, assuming that $\nu$ is a geometric sequence is extremely restrictive. Instead of directly using geometric sequences, we use bijective deformations of positive convergent geometric sequences as the upper bounds of $\nu$.
Let $\mathbb{I}\left(\mathbb{R}\right)$ be the set of all intervals on $\mathbb{R}$ with nonempty interiors. We define, for a function $f$ on reals, the set of intervals where $f$ is strictly increasing and continuous that is,
[…]
If $I\in\mathrm{SIC}(f)$, $f$ is bijective from $I$ to $f(I)$ and we denote by ${{f}_{\left(I\right)}^{-1}}$ the inverse of $f$ on $I$. Note that ${{f}_{\left(I\right)}^{-1}}$ is also strictly increasing and continuous from $f(I)$ to $I$. Finally, we pay particular attention to functions that are strictly increasing and continuous on $[0,1]$ and we write
A convergent geometric sequence is required to complete the upper bound on $\nu$ and we introduce
$\mathbb{\Gamma}(\nu):=\{(h,\beta)\in\Omega([0,1])\times(0,1):\nu_{k}\leq h(\beta^{k}),\ \forall\,k\in\mathbb{N}\}$
and
$\mathbb{\Gamma}_{\rm fun}(\nu):=\{h\in\Omega([0,1]):\exists\,\beta\in(0,1)\text{ s.t. }(h,\beta)\in\mathbb{\Gamma}(\nu)\}.$
Finally, we introduce for $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$:
$\mathbb{\Gamma}_{\rm sc}(\nu,h):=\{\beta\in(0,1):(h,\beta)\in\mathbb{\Gamma}(\nu)\}.$
From $(h,\beta)\in\mathbb{\Gamma}(\nu)$, we can define
$\mathfrak{F}_{\nu}(k,h,\beta):=\dfrac{\ln\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\nu_{k})}{\ln(\beta)}\text{ if }\nu_{k}>h(0)$ (4)
for which we have
$j\geq\lfloor\mathfrak{F}_{\nu}(k,h,\beta)\rfloor\implies\nu_{j}<\nu_{k}$
The point requiring special consideration is the strict positivity of ${{h}_{\left([0,1]\right)}^{-1}}(\nu_{k})$ or, equivalently, the fact that $\nu_{k}>h(0)$. This depends on the choice of $h$ and we define the notion of usefulness.
Definition 1 (Useful Strictly Increasing Continuous Functions).
Suppose that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$. A function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ is said to be useful for $\nu$ if the set
$\mathcal{S}(\nu,h):=\{k\in\mathbb{N}:\nu_{k}>h(0)\}$ (5)
is nonempty.
By extension, a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$ if $h$ is useful for $\nu$.
Remark 1.
Note that as $h\in\mathbb{\Gamma}(\nu)$ is continuous, we have $\limsup_{k\to+\infty}\nu_{k}\leq h(0)$ and thus $\mathcal{S}(\nu,h)\subset{\mathrm{R}}_{\nu}^{s}$.
One of the main results of [Adj25b] is the existence of a useful element in $\mathbb{\Gamma}(\nu)$ when ${\mathrm{R}}_{\nu}^{s}$ is nonempty.
Theorem 1.
The following statements are true.
@@
    $\nu_{\rm opt}<+\infty$ if and only if $\mathbb{\Gamma}(\nu)\neq\emptyset$;
@@
    ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if there exists $(h,\beta)\in\mathbb{\Gamma}_{\rm fun}(\nu)$ that is useful for $u$.
2.1.3 Problem statement
Even if we have proved that $\mathbb{\Gamma}(\nu)$ is nonempty, the proof furnished in [Adj25b] exhibits an affine function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ for which the slope and y-intercept depend on $\mathbf{K}^{s}_{\nu}$ and $\nu_{\rm opt}$. This leads to considering alternative approaches to construct an element in $\mathbb{\Gamma}(\nu)$.
Problem 1.
In [Adj25a], we proposed to build $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a $\mathcal{K}\mathcal{L}$ stability certificate or a Lyapunov function. Problems of the form (DPCP) can have a finite optimal value and are such that ${\mathrm{R}}_{\nu}^{s}$ is nonempty without admitting a classical $\mathcal{K}\mathcal{L}$ stability certificate or a classical Lyapunov function. This is the case for the example developed in Section 3. To go further Problem 1, we are interested in a full characterization of the elements of $\mathbb{\Gamma}(\nu)$ that subsumes the previous results based on classical stability tools. Thus, the main contribution of this paper is the characterization of the elements of $\mathbb{\Gamma}(\nu)$ from an extension of $\mathcal{K}\mathcal{L}$ and the Lyapunov functions. We will prove that solving Problem 1 is equivalent to constructing a weaker version of $\mathcal{K}\mathcal{L}$ certificate and a weaker version of the Lyapunov functions. We will prove that we can construct an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from any extended $\mathcal{K}\mathcal{L}$ certificate and that an extended $\mathcal{K}\mathcal{L}$ certificate can be constructed from any $(h,\beta)\in\mathbb{\Gamma}(\nu)$. The same results hold for our extended class of Lyapunov functions.
2.2 Classical $\mathcal{K}\mathcal{L}$ and Lyapunov approaches
The use of $\mathcal{K}\mathcal{L}$ and Lyapunov functions outside the scope of stability theory seems to be atypical and we bring some intuitions about the use in the context of solving Problem 1. In fact, classical $\mathcal{K}\mathcal{L}$ and Lyapunov functions furnish more or less naturally a function in $\Omega([0,1])$ and $\beta\in(0,1)$. To obtain an element of $\mathbb{\Gamma}(\nu)$ from a $\mathcal{K}\mathcal{L}$ certificate or a Lyapunov function requires a restrictive inequality or the continuity of $\varphi$. From the positivity of $\mathcal{K}\mathcal{L}$ and Lyapunov functions, usefulness holds if and only if the sequence $\nu$ has a positive term. All the details on the approach can be found in [Adj25a].
2.2.1 The classical $\mathcal{K}\mathcal{L}$ approach
First, recall that a $\mathcal{K}\mathcal{L}$ function $\gamma:\mathbb{R}_{+}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ is such that
@@ 2.2.1 The classical $\mathcal{K}\mathcal{L}$ approach
    For all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)\in\mathcal{K}_{\infty}$ i.e. the set of strictly increasing continuous functions $\mathbb{R}_{+}\to\mathbb{R}_{+}$ that are zero at zero and whose the limit at $\infty$ is $\infty$;
@@
    For all $s\in\mathbb{R}_{+}$, $t\mapsto\gamma(s,t)\in\mathcal{L}$ i.e. the set of decreasing continuous functions $\mathbb{R}_{+}\to\mathbb{R}_{+}$ whose the limit at $\infty$ is 0;
The dynamical system $(X^{\mathrm{in}},T)$ admits a $\mathcal{K}\mathcal{L}$ certificate if and only if there exists $\gamma\in\mathcal{K}\mathcal{L}$ such that for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$
$\|T^{k}(x)\|\leq\gamma(\|x\|,k)$ (6)
If the inequality (6) holds, the famous result of Sontag [Son98, Proposition 7] directly provides a potential pair $(h,\beta)$. Recall that the result asserts that for any $\gamma\in\mathcal{K}\mathcal{L}$ there exists $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ verifying $\gamma(s,t)\leq\theta_{1}(\theta(s)e^{-t})$ for all $(s,t)\in\mathbb{R}_{+}\times\mathbb{R}_{+}$. Then, if $X^{\mathrm{in}}$ is bounded, we define
$h:s\to\theta_{1}(\theta_{2}(s\sup_{x\in X^{\mathrm{in}}}\|x\|))\in\Omega([0,1]).$ (7)
The difficulty is to exploit the inequality (6) to obtain $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$. This requires an additional assumption on $\varphi$. For example, if $\varphi$ is continuous, we obtain the desired upper bound from the famous Hahn/Khalil result [Hah67] and [Kha02, Lemma 4.3/Appendix C.4]. This result asserts that for any positive definite222a function is positive definite if it is strictly positive except at 0 where it is null continuous coercive333a function $f$ is said to be coercive (or radially unbounded) if $\lim_{\|x\|\to+\infty}f(x)=+\infty$ function $f$ there exist $\overline{\alpha_{1}},\overline{\alpha_{2}}\in\mathcal{K}_{\infty}$ satisfying:
$\overline{\alpha_{1}}\circ\|\cdot\|\leq f\leq\overline{\alpha_{2}}\circ\|\cdot\|.$ (8)
From Assumption 1 $\max\{\varphi,\|\cdot\|\}$ is positive definite, coercive, and continuous if $\varphi$ is. Hence
$(\alpha_{\varphi}\circ h,e^{-1})\in\mathbb{\Gamma}(\nu)\text{ where }\alpha_{\varphi}\in\mathcal{K}_{\infty}\text{ is such that }\varphi\leq\max\{\varphi,\|\cdot\|\}\leq\alpha_{\varphi}\circ\|\cdot\|.$ (9)
We can relax the continuity assumption on $\varphi$ and instead imposing that, for some $\psi:\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\sup_{x\in X^{\mathrm{in}}}\psi(x)<+\infty$
$\varphi(T^{k}(x))\leq\gamma(\psi(x),k),\text{ for all }x\in X^{\mathrm{in}}\text{ and }k\in\mathbb{N}.$ (10)
This consists in replacing in the inequality (6) the norm on the left-hand side by $\varphi$ and the norm on the right-hand side by a nonnegative function $\psi$ with a strictly positive supremum on $X^{\mathrm{in}}$. In this case
$(s\mapsto\theta_{1}(\theta_{2}(s\sup_{x\in X^{\mathrm{in}}}\psi(x))),e^{-1})\in\mathbb{\Gamma}(\nu).$ (11)
Finally, as stated above, usefulness in the context of $\mathcal{K}\mathcal{L}$ certificate is equivalent to the existence of a strictly positive term for $\nu$.
2.2.2 The Lyapunov approach
In [Adj25c], we consider Lyapunov functions to solve Problem (DPCP). In [Adj25c], the dynamical system was a stable affine system, and the objective function $\varphi$ was a quadratic function. Recently, in [Adj25a], we generalized the approach for all discrete-time systems that admit a Lyapunov function. Recall that a function $V:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is a (global) Lyapunov function if and only if
$\alpha_{1}\circ\|\cdot\|\leq V\leq\alpha_{2}\circ\|\cdot\|\text{ for some }\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ (12)